The labor relations manager of a large corporation wished to study the absenteeism among workers


QUESTION 3 

The labor relations manager of a large corporation wished to study the absenteeism among workers at the company's central office during the last year. a random sample of 25 workers revealed the following: an average of 9.7 days; a standard deviation of 4.0 days; and 12 employees were absent more than 10 days. In estimating a 95% confidence interval estimate of the average number of days absent for the company's workers last year what degrees of freedom will you use in calculating the margin of error? 


QUESTION 4 

The labor relations manager of a large corporation wished to study the absenteeism among workers at the company's central office during the last year. a random sample of 25 workers revealed the following: an average of 9.7 days; a standard deviation of 4.0 days; and 12 employees were absent more than 10 days. ln estimating a 95% confidence mental estimate of the average number of days absent for the company's workers last year what will be the value of your t statistic used in calculating your margin of error? 


QUESTION 5 

The labor relations manager of a large corporation wished to study the absenteeism among workers at the company's central office during the past year a random sample of 25 workers revealed the following: an average of 9.7 days; a standard deviation of 4.0 days; and 12 employees were absent more than 10 days in estimating a 95% confidence interval estimate of the average number of days absent for the co-pays workers last year what is your margin of error? 


QUESTION 6 

The labor relations manager of a large corporation wished to study the absenteeism among workers at the company's central office during the last year a random sample of 25 workers revealed the following: an average of 9.7 days; a standard deviation of 4.0 days; and 12 employees were absent more than 10 days. In estimating a 95% confidence interval estimate of the average number of days absent Of the company's workers last year what would be the value of your UPPER limit? 
Answer #1

3)

df = n -1 = 25-1 = 24
4)

for 95 %
t- critical value = =T.INV(0.975,24)
= 2.063899

5)
margin of error = t * sd/sqrt(n)
= 2.063899 * 4/ sqrt(25)
= 1.6511192

6)
upper limit = mean + margin of error
= 9.7 + 1.6511192
= 11.3511192